OpenStudy (anonymous):

How does 8z-z/2=z/2+15 turn into: 15z/2=15? I understand taking away z/2, but I don't understand where the 15z came from.

OpenStudy (campbell_st):

well 8 - 1/2 = 7 1/2 or 15/2

OpenStudy (anonymous):

mulitply both sides by 2 you get 16z-z=z+30 combined terms 15z=z+30 -z -z 14z=30 probably easier

OpenStudy (anonymous):

so it wouldnt turn into 15z/2=15 anyway :3

OpenStudy (anonymous):

would be 14z/2=15

OpenStudy (campbell_st):

well its 15z/2 = z/2 + 15 subtract z/2 from both sides 14z/2 = 15 or 7z = 15 z = 15/7